OpenStudy (swissgirl):

Give a counterexample for the following statement: \( \overline{\overline A } \leq \overline{\overline B } \text{ implies } A \subset B \)

OpenStudy (anonymous):

What does \(\overline{\overline{S}}\) mean? Is that the cardinality of the set?

OpenStudy (swissgirl):

yes

OpenStudy (anonymous):

\(A=\mathbb{Z}\) \(B=\mathbb{R}\times \mathbb{R}\)

OpenStudy (anonymous):

There are numerous counterexamples.

OpenStudy (swissgirl):

Well \( \mathbb{Z}\) is a subset of \( \mathbb{R} \)

OpenStudy (anonymous):

The cartesian product of two sets is a set of ordered pairs.

OpenStudy (swissgirl):

Can you explain to me what cardinality actually means

OpenStudy (turingtest):

how about B is all even numbers and A is the first ten odd number? (I don't know if you need infinite sets or what)

OpenStudy (swissgirl):

not sure either. i am just trying to figure out all this stuff

OpenStudy (swissgirl):

but I am assuming that it refers to denumerable sets

OpenStudy (anonymous):

No, uncountable sets still have cardinalities.

OpenStudy (turingtest):

I have never taken this kind of math, but cardinality as I basically understand it is how many elements it has (I don't know about sets of ordered pairs)

OpenStudy (anonymous):

Ordered pairs act as elements of the set. They do not change anything.

OpenStudy (turingtest):

so what's the matter with my finite example?

OpenStudy (anonymous):

Nothing!

OpenStudy (turingtest):

oh, sweet :)

OpenStudy (swissgirl):

ok so how abt this example: A= \( \mathbb{Q^-} \) B= \( \mathbb{Q^+} \)

OpenStudy (swissgirl):

hahahahahah

OpenStudy (swissgirl):

like do the cardinalities equal each other?

OpenStudy (anonymous):

Yes.

OpenStudy (turingtest):

they are both\[\aleph_0\]right? (side question)

OpenStudy (swissgirl):

lol I think soooo

OpenStudy (anonymous):

Yep. They can both be put on a one-to-one correspondence with the positive integers.

OpenStudy (turingtest):

I gotta take this formally some day... thanks!

OpenStudy (swissgirl):

THANKS GUYSSSS :))